Systems and Network Management

Configuring Cricket to Monitor SNMP Objects

Contents

1 Procedure  2
1.1 Open the Documentation . . . . . . . . . . . . . . . . . . . . . . . . .  2
1.2 Installing the Necessary Perl modules . . . . . . . . . . . . . . . . . .  2
1.2.1 Installing the Required Perl Modules from CPAN . . . . . . .  2
1.2.2 Installing the other two Perl Modules . . . . . . . . . . . . . .  4
1.3 Creating the cricket account . . . . . . . . . . . . . . . . . . . . . .  4
1.4 Unpacking Cricket and running configure . . . . . . . . . . . . . . .  5
1.5 Copy the Config Tree and Modify it . . . . . . . . . . . . . . . . . . .  6
1.6 Adding a cron entry for Cricket . . . . . . . . . . . . . . . . . . . . .  6
1.7 Setting up the public html directory in cricket’s home directory for Apache  7
1.8 Configuring the Apache Web Server . . . . . . . . . . . . . . . . . . .  7
2 Adding a new set of Graphs and Targets for Cricket to Monitor  9
2.1 Cricket’s Configuration Tree . . . . . . . . . . . . . . . . . . . . . . .  10
2.1.1 A Closer Look at the Config Tree . . . . . . . . . . . . . . . .  10
2.2 The servers subtree . . . . . . . . . . . . . . . . . . . . . . . . . . . .  11
2.2.1 Variable Expansion . . . . . . . . . . . . . . . . . . . . . . . .  11
2.3 Tables and Instance Numbers . . . . . . . . . . . . . . . . . . . . . .  11
2.3.1 The listInterfaces Utility . . . . . . . . . . . . . . . . . . .  13
2.4 Monitoring High-End Cisco Switches: the genCatConfig Utility …  13
2.5 Monitoring 35xx Cisco Switches and Routers . . . . . . . . . . . . . .  14
2.6 Running the Name Service Caching Daemon . . . . . . . . . . . . . .  14
2.7 Installing the Red Hat Updates . . . . . . . . . . . . . . . . . . . . .  15

Background:

There are two free software packages for monitoring snmp objects from a web interface.

The most widely known is mrtg, the “Multi Router Traffic Grapher.” mrtg is well suited to monitoring network traffic from router interfaces, but has certain limitations in many other situations. Red Hat Linux comes with mrtg as a ready-built software package, and it is very easy to set it up to monitor network traffic. See http://people.ee.ethz.ch/~oetiker/webtools/mrtg/.

The newer package is called Cricket, and this is what we will install and configure today. Cricket has greater flexibility, and can be extended in more ways. Cricket is

available from http://cricket.sourceforge.net/.

Both packages are written mostly in Perl. There are links to both of these software packages from the subject web site.

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

2

1 Procedure

There are eight steps to install and configure Cricket, as I demonstrated at the last lecture: • install the necessary Perl modules using cpan

• create a user called cricket

• unpack the Cricket tarball and run the configure script

• copy the sample configuration files into a directory called cricket-config

• modify the configuration to point to what you want to monitor

• configure cron to run the data collector

• create a public html directory, and set up the programs in that directory

• configure Apache so that it will run the programs

1.1 Open the Documentation

1. Go to the Cricket installation web page at http://cricket.sourceforge.net/support/doc/beginner.html.

You can get there from the main Cricket page by following these links: support → documentation → Installing Cricket for the Complete Beginner.

1.2 Installing the Necessary Perl modules

There are several Perl modules that Cricket requires. Some of these are already installed on your computer, some are not. You will install them with the cpan program, which downloads them from the Internet, compiles, tests, then installs them if they pass all tests.

The modules to install from cpan include:

• MD5

• DB File

• Date::Parse

• Time::HiRes

• CGI

• Net::SNMP 1.2.1 Installing the Required Perl Modules from CPAN

1. First, you need the cpan Perl module. People who cloned their hard disk from mine will find that you need to install it:

  $ sudo rpm -Uhv \
    /home/nfs/rh-8.0-updated/RedHat/RPMS/perl-CPAN-1.61-55.i386.rpm

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

3

As usual, use the [figure: tab]key to complete the file name; I don’t remember these version numbers myself, of course!

2. Secondly, you need the development software package for the Berkeley Database system required to build DB File:

$ sudo rpm -Uhv \
/home/nfs/rh-8.0-updated/RedHat/RPMS/db4-devel-4.0.14-14.i386.rpm

3. Start the cpan modules’ shell method from your own (not cricket’s) account:

$ sudo perl -MCPAN -e shell

4. The shell will begin its interactive configuration process. Accept all defaults by pressing [figure: Enter], except for the following:

  • When prompted “CPAN build and cache directory?” enter: /var/cache/cpan

Do not accept the default in your home directory; it will not work, since the root user has no right to write to your nfs mounted home directory.

• When asked to choose a Policy on building prerequisites, choose “follow”:

    Policy on building prerequisites (follow, ask or ignore)? [ask] follow

Some Perl modules depend on others. When installing a module that depends on others, a selection of “ask” here means the cpan shell will prompt you and require you to press [figure: Enter]to install the prerequisite module. The selection “follow” means, “Don’t ask me; just go ahead, download and install any modules that are needed to install this one”.

  • For “Your http proxy?” enter http://hqproxy.vtc.edu.hk:8080/

• When asked to pick a nearby continent, enter 5 to select (5) North America

    (Do not select (2) Asia, 2 ftp://ftp.pacific.net.hk/pub/mirror/CPAN/,

because like many Hong Kong mirror sites, it is not maintained properly)

  • When asked to “Select your country”, enter 3 to select (3) United States.

• When asked “Select as many URLs as you like”, I suggest just enter the first few numbers (corresponding to the first few urls); I put “1 2 3 4 5 6 7”. The numbers are just separated by spaces.

• Configuration should complete.

5. Now, at the cpan> prompt, type:

cpan> install MD5
cpan> install Date::Parse
...

and so on, installing all the six modules listed above.

6. For more information about the cpan software, type “help”. There is a manual page; type perldoc CPAN, or man CPAN to read it.

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

4

1.2.2 Installing the other two Perl Modules

The other two Perl modules are not available from cpan; see the Cricket documentation for the location. We start here assuming that ictlab is mounted on /home/nfs. I have set up appropriate automounter entries in our ldap directory server so that when you change to or list this directory, it will mount the network file system by nfs.

Do this in your own account.

Installing SNMP Session:

1. Unpack the SNMP Session tarball into your (not cricket’s) home directory:

     $ cd
     $ tar xvzf /home/nfs/snmp/SNMP_Session-0.94.tar.gz
     $ cd SNMP_Session-0.94
     $ perl Makefile.PL
     $ make
     $ sudo make install

The SNMP Session Perl modules are now installed.

          You can get the latest copy of this module from ftp://ftp.switch.ch/software/sources/network/snmp/perl/.
          The web page is at http://www.switch.ch/misc/leinen/snmp/perl/.

Installing the Round Robin Database: This is available as two rpm packages, built

from source rpms downloaded from http://rpmfind.net/linux/falsehope/home/gomez/rrdtool/rrdtool-1.0.39-1.7.2.src.rpm.

1. Simply install using the rpm program:

     $  sudo rpm -Uhv \
     /home/nfs/redhat/contrib/rrdtool-*1.0.39-1.8.0.i386.rpm --nodeps

After this, all the required Perl packages are installed.

1.3 Creating the cricket account

Here you create a local user. You will install the software into the home directory of this user.

Setting your path

1. Look at your path by typing:

     $ echo $PATH

2. If the directories /sbin and /usr/sbin are not already on your path, add them:

(a) edit your login script:

         $  emacs ∼/.bash_profile &

(b) Add a line like this at the end of your login script:

         export PATH=$PATH:/sbin:/usr/sbin

(c) Copy and paste this line into your terminal window.

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

5

Now create the cricket account:

1. Make sure that your useradd program is configured to create local accounts (see the handout I wrote about this for more information):

$ sudo mkdir /home2
$ sudo useradd -D -b /home2

2. Create a local user cricket:

$ sudo useradd -c "Cricket Manager" cricket

3. Give this user a password:

$ sudo passwd cricket

1.4 Unpacking Cricket and running configure

The Cricket tarball is available for download either from the Cricket web site, or you can install it directly from our server:

1. Open a window and become cricket:

$ xhost +localhost
$ su - cricket

The first command lets other users besides you (such as cricket) display on the local X server. The minus sign ‘-’ runs cricket’s login scripts.

2. Change to cricket’s home directory and unpack the Cricket tarball:

$ cd
$ tar xvzf /home/nfs/snmp/cricket-1.0.3.tar.gz

3. Now create a symbolic link to the Cricket directory:

$ ln -s cricket-1.0.3 cricket

4. Now run the configure script, which puts the path to perl in the first line of the Perl programs:

$ cd ∼/cricket
$ ./configure
5. Now copy the file cricket-conf.pl.sample to cricket-conf.pl:
$ cd ∼/cricket
$ cp cricket-conf.pl.sample cricket-conf.pl

6. Edit the file ∼/cricket/cricket-conf.pl, and change the line:

$gCricketHome = "/home/cricket";

to

$gCricketHome = "/home2/cricket";

Nick Urbanik <nicku(at)nicku.org>  ver. 1.19

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

6

1.5 Copy the Config Tree and Modify it

Cricket uses a directory of configuration files that the authors call a config tree. Here we copy the sample config tree, and modify it.

1. First copy the sample tree as the user cricket:

$ cd
$ cp -a cricket/sample-config cricket-config

2. Now copy the configuration that I wrote for the demonstration in the lecture theatre. Here I assume that ictlab is mounted on /home/nfs.

$ cd ∼/cricket-config
$ cp -a /home/nfs/snmp/servers .
#      This is a dot __________^

There are two files in this directory: Defaults and Targets. These tell Cricket to monitor the system load on our server ictlab, and also the amount of free swap space, and the amount of free memory. You will later need to modify the config tree to monitor other devices.

    If you copied the two files from /home/nfs/snmp/servers into ∼/cricket-config

instead of into the directory ∼/cricket-config/servers, then the collector will not work. If you have done this, simply go back and copy the sample configuration files back again.

3. Every time you modify the config tree, you need to run the compile script (as the

user cricket):
$ ∼/cricket/compile

4. Cricket collects data from data sources using a script called the collector. This will be run by cron. First, check that the config tree has been set up correctly:

$ ∼/cricket/collector /servers

This should run without errors, and five data items should be shown: the three values for system load, and the free swap space in kilobytes, and free memory.

1.6 Adding a cron entry for Cricket

Cricket collects the information it graphs using cron, every five minutes. To use cron to gather the information for Cricket, you need to add an entry to cricket’s crontab like this. First, you need to be cricket.

1. The cron program collects information about the network periodically using script called collect-subtrees. This script reads a configuration file ∼cricket/cricket/ subtree-sets. Edit this file:

$ emacs ∼/cricket/subtree-sets &

2. Add a new set to it like this:

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

7

     set linuxservers:
             /servers

The script collect-subtrees takes one parameter, the set of data to collect. Here our set is called linuxservers so we use that parameter in our cron entry.

3. Start editing cricket’s crontab:

     $ crontab -e

4. In the editor, you need to add an entry like this:

     */5 * * * * $HOME/cricket/collect-subtrees linuxservers
     which will run the program $HOME/cricket/collect-subtrees linuxservers

every five minutes.

1.7 Setting up the public html directory in cricket’s home directory for Apache

I found it necessary to make a minor change from the instructions for putting the cgi programs in the public html directory. This is what I did, as the user cricket:

$ cd
$ mkdir -p public_html/cricket
$ cd public_html/cricket
$ ln -s ∼/cricket/VERSION
$ ln ∼/cricket/grapher.cgi .
$ ln ∼/cricket/mini-graph.cgi .
$ ln -s ∼/cricket/lib ∼/cricket/images .
# Note: this is a dot ------------------^

The only difference is that I made a hard link, rather than a symbolic link, to the cgi programs.

I have also added a link to the documentation, so that you can easily read it online:

$ ln -s ∼/cricket/doc ∼/public_html

You can read the Cricket documentation at http://localhost/~cricket/doc/ after you have set up Apache—see the next section.

1.8 Configuring the Apache Web Server

Now configure Apache to run Cricket:

1. Edit Apache’s configuration file:

     $ sudo -v
     $ sudo emacs /etc/httpd/conf/httpd.conf &

2. Search for the string “UserDir” in your editor

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

8

3. Comment out the line

    UserDir disable

by putting a hash “#” in front of it:

    #UserDir disable

and uncomment the line

    UserDir public_html

by removing the hash character.

4. Search for the “Directory” section for public html.

5. If the section is commented out, copy it and uncomment the copy.

6. You need to add the option ExecCGI, and should end up with something like this:

<Directory /home2/*/public_html>
   Options MultiViews Indexes SymLinksIfOwnerMatch ExecCGI
</Directory>

Note that the manual for Apache is included on your hard disk. When Apache is running, you will find it at this location: http://localhost/manual/. The manual is very complete, and explains every possible configuration option of Apache.

7. You also need to uncomment the line in Apache’s configuration file:

AddHandler cgi-script .cgi

8. Note too that cricket’s home directory, and ∼cricket/public html, and ∼cricket /public html/cricket all need to have the group write permission removed, but the execute permission added for group and others. This is to satisfy the requirements of suexec, which is the mechanism by which Apache executes Cricket’s programs as the user cricket.

The complete documentation for suexec is included with the online Apache documentation on your hard disk. Okay, it seems nobody understands that, so here’s the recipe:

    $ chmod go=x ∼
    $ chmod go=rx ∼/cricket ∼/public_html ∼/public_html/cricket

The permissions should look something like this:

      $ ls -ldL ∼ ∼/cricket ∼/public_html ∼/public_html/cricket
      drwx--x--x    7 cricket  cricket      4096 Dec 18 13:44 /home2/cricket
      drwxr-xr-x    7 cricket  cricket      4096 Dec 11 17:21 /home2/cricket/cricket
      drwxr-xr-x    2 cricket  cricket      4096 Dec 11 17:55 /home2/cricket/public_html
      drwxr-xr-x    2 cricket  cricket      4096 Dec 11 17:55 /home2/cricket/public_html/cricket

You could remove write permission for group from all the files like this:

Nick Urbanik <nicku(at)nicku.org>  ver. 1.19

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

9

$ chmod -R g-w ∼

which will remove the write permission for group of all files below ∼/public html.

9. You can use the serviceconf program to start Apache’s httpd service in runlevels 3, 4 and 5.

Alternatively you can use the same method to start and maintain the Apache service as with nscd given on page 14.

$ sudo service httpd start
$ sudo chkconfig --level 345 httpd on

Testing, and using the log files:

1. Now open a web browser, and go to http://localhost/~cricket/cricket/grapher.cgi

2. If you see a server error message:

(a) View the log files for Apache to see what is happening. I suggest change the group ownership of the log file directory to you, so that you can see the files without being root:

$ ls -ld /var/log/httpd
drwx------    2 root     root         4096 Oct 31 17:52 /var/log/httpd
$ sudo chgrp nicku /var/log/httpd
$ sudo chmod g+rx /var/log/httpd
$ ls -ld /var/log/httpd
drwxr-x---    2 root     nicku        4096 Oct 31 17:52 /var/log/httpd

Now change to the web server’s log directory and examine the logs:

$ cd /var/log/httpd
$ ls -ltr
$ tail -f error_log

(b) You may also want to open two more windows, and leave them open, running

tail -f on suexec log and access log. In particular, monitoring suexec log

will let you know about permission problems when accessing the programs.

(c) After modifying Apache’s configuration, tell Apache to reload it:

$ sudo service httpd graceful

3. If you see the Cricket page, click on servers, then load, and you should see the graphs of system load there.

2 Adding a new set of Graphs and Targets for Cricket to Monitor

Now this is where you use your knowledge of snmp. Refer to the page Setting up New De-

vices in Cricket, at http://cricket.sourceforge.net/support/doc/new-devices.html,
or http://localhost/~cricket/doc/new-devices.html if you make the link I described

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

10

at the end of section 1.7 on page 7. There is a useful set of configurations in http://www.gnac.com/techinfo/cricket_contrib/index.html.

The reference for the configuration of Cricket is at http://cricket.sourceforge.net/support/doc/reference.html.

Cricket has a very active mailing list for users, about ten posts per day. I had an perfect solution to my problem in less than two hours. The subscription information is

available at http://cricket.sourceforge.net/support/.

2.1 Cricket’s Configuration Tree

The directory ∼/cricket-config contains a directory tree. This directory tree contains the configuration of cricket. It is called a “config tree” because the configuration files at the higher levels provide default values, and some of these values are overridden by the configuration files at the lower levels.

Every directory in ∼/cricket-config contains a file called Defaults.

The top level configuration file, ∼/cricket-config/Defaults, contains a very large number of useful definitions. In particular, it is set up so that all you need to plot network traffic is in place. The oids are provided for all the useful items in the Mib-2 Interfaces table. Look at this file; you will see that the default graphs are graphs of network traffic in and out of a network interface.

Each chunk of the file begins with a word such as “target” or “OID”. After that word, they are different, but you will see many things such as

rrd-datafile = %dataDir%/%auto-target-name%.rrd.

The percent signs indicate a variable that is expanded somewhere else in the configuration tree. Many of these variables have reasonable defaults that are defined at the top level of the tree. You only need to fill in the values for some (such as the hostname of the target) further down in the config tree.

2.1.1 A Closer Look at the Config Tree

The configuration is divided into dictionaries. Below some dictionaries, there are tags and values, arranged like this:

dictionary − name tag1 = value1 tag2 = value2

    ...

These are the dictionaries available:

target A device that we monitor is a target. There is one rrd file for each target.

datasource A Data source is one line on a graph. One rrd can have many datasources in it. All datasources in one rrd need to be updated by the collector at once.

targetType Determines what kinds of data sources there are for the target with this targetType. The collector uses the targetType to decide what data to fetch, and how to fetch it.

graph Determines how individual data sources are graphed.

color Maps colour names to html colours. Probably no need to change this.

oid Maps an oid to a name

html determines what html goes into each page

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

11

event draws vertical bars on the graph to indicate events, such as the end of one day and the start of the next. No need to change this.

rra Holds parameters used by Cricket when creating a new rrd file. No need to change this.

2.2 The servers subtree

Let us consider how the configuration for /servers works I decided to monitor the system load average and the available free memory. There are five data sources here, three for the load averages, one for available swap,

and one for free memory. These are listed under the targetType dictionary name.

2.2.1 Variable Expansion

When the collector gets the system load data from ictlab, it uses an snmp url something like this:

   snmp://public@nicku.org:161:2.0:5:1.0:1/1.3.6.1.4.1.2021.10.1.3.1

But where did all that come from?

Let’s look in the Target dictionary in the top level ∼/cricket-config/Defaults file:

Target    --default--
    snmp-host           = %auto-target-name%
    snmp-community      = public
    snmp-port           = 161
    snmp-timeout        = 2.0
    snmp-retries        = 5
    snmp-backoff        = 1.0
    snmp-version        = 1
    snmp                = %snmp-community%@%snmp-host%:%snmp-port%:
%snmp-timeout%:%snmp-retries%:%snmp-backoff%:%snmp-version%

So what about %auto-target-name%? This is automatically set by Cricket to the name of the target. So let’s see part of the file ∼/cricket-config/servers/Targets:

target  nicku.org
    target-type =   linux-machine

Finally, let’s see how the oid and the data source is defined in ∼/cricket-config

/servers/Defaults:
OID     laLoad1min  1.3.6.1.4.1.2021.10.1.3.1
datasource  laLoad1min   ds-source   =   snmp://%snmp%/laLoad1min

So default values mostly come from the top level of the configuration tree and are inherited below. You can override any default defined at the top level with values you define further down.

2.3 Tables and Instance Numbers

We examined the Interfaces table of Mib-2 in some detail. Each row in the table corresponds to one network interface. How do we use snmp to get information on a particular network interface? Let’s look at ifInOctets, in the mib file /usr/share/snmp/mibs

/RFC1213-MIB.txt. The definition of ifInOctets is:

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

12

ifInOctets OBJECT-TYPE
    SYNTAX  Counter
    ACCESS  read-only
    STATUS  mandatory
    DESCRIPTION
            "The total number of octets received on the
            interface, including framing characters."
    ::=  ifEntry 10

Let’s get the oid of this table entry:

$ snmptranslate -IR -Of ifInOctets
.1.3.6.1.2.1.2.2.1.10

Good, now let’s see the value on ictlab:

$ snmpwalk -c public ictlab .1.3.6.1.2.1.2.2.1.10
interfaces.ifTable.ifEntry.ifInOctets.1 = Counter32: 752366405
interfaces.ifTable.ifEntry.ifInOctets.2 = Counter32: 791891118

Hey, there are two of them! What are they? Well, Mib-2 provides a description of the interface in ifDescr; let’s see what that says:

$ snmpwalk -c public ictlab ifDescr
interfaces.ifTable.ifEntry.ifDescr.1 = lo
interfaces.ifTable.ifEntry.ifDescr.2 = eth0

Okay, so the entries in the ifTable with a .1 at the end on ictlab relate to the loopback interface, while the entries with a .2 at the end are about the Ethernet interface.

This number at the end of the oid is called an instance number, and is the way that snmp selects items from a table. Notice that the instance numbers begin at 1 for table entries. What about for scalars?

$ snmpwalk -c public ictlab sysUpTime
system.sysUpTime.0 = Timeticks: (103614809) 11 days, 23:49:08.09

The instance number for a scalar is always zero.

Typically, the loopback interface is the first entry in the table, and other interfaces follow. However, snmp does not prevent a router from changing the order of the interfaces. This could be a real problem; after rebooting a router, you may find that the graph for your subnet is now collecting data for a different subnet! Luckily, Cricket has a solution to this

called instance mapping, described in http://cricket.sourceforge.net/support/doc/inst-mapping.html.

It is also possible to directly provide a value for the instance number for a target to graph, instead of using instance mapping. For example, you could set

target hostname-of-server-or-router
    inst = 2

to draw graphs of interface two of the machine called hostname-of-server-or-router.

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

13

2.3.1 The listInterfaces Utility

A very nice Cricket utility is provided to automatically find the names of all the interfaces on a server or router, then map each name to an instance number. This is described

in http://cricket.sourceforge.net/support/doc/beginner.html. The name of the

utility is listInterfaces. If main-router is the hostname of a server or router that you want to plot statistics of the network interfaces from, you can do:

$ ∼/cricket/util/listInterfaces main-router > interfaces
The file interfaces can then be used in the ∼/cricket/cricket-config/router-interfaces

directory to specify how to plot each interface. All that is left is simply to put an entry in subtree-sets so that the collector will collect the data from the server or router, and then run the compile script. Cricket provides enormous flexibility once it is set up.

2.4 Monitoring High-End Cisco Switches: the genCatConfig Utility

To monitor high-end Cisco Catalyst^{TM} switches, such as our 6509 (but not Cisco Catalyst 3500XL switches), there is a tool available for automatically generating the configuration

for Cricket. It is currently available from http://www.certaintysolutions.com/tech-advice/cricket-contrib/;

you can download the program from that page. Note that all these programs are available

from our own server ictlab, by many protocols, including http://nicku.org/ftp/snmp/cricket-contrib,
and by nfs from ictlab:/var/ftp/pub/snmp.

Read the file ∼/cricket/util/README.genCatConfig for the full documentation; this is just a short summary. To install it,

• become cricket, and

• untar the package into /tmp. Then

   • copy the directory /tmp/sample-config/catalysts and its contents into ∼/cricket-config.

• Copy the contents of /tmp/util into ∼/cricket/util.

• Edit the two executables and change the first line in each file from

     #!/usr/local/bin/perl

to

     #! /usr/bin/perl

• Make a host table entry for the router in /etc/hosts.

   • Change directory to ∼/cricket-config/catalysts

• Then run

$ ∼/cricket/util/genCatConfig -2 -v -C ⟨community-string⟩ ⟨router-hostname⟩

replacing ⟨router-hostname⟩ with the host name of the Catalyst router that you put into the /etc/hosts file, and ⟨community-string⟩ with its read-only community string.

   • Add /catalysts/router-hostname to Cricket’s subtree-sets file.

• Run ∼/cricket/compile. It should now collect data. Test it using the collector.

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

14

2.5 Monitoring 35xx Cisco Switches and Routers

According to the genCatConfig documentation:

Some Catalysts (e.g. 2900XL and 3500XL switches) run router ios and respond to a different set of mibs. They should be handled as routers, not switches. (see genRtrConfig).

So download genRtrConfig from the same place as genCatConfig above. The installation process is similar, but not identical. In particular, the Cisco 3500XL series does not support the 64 bit counters, so we have to use snmp version 1! Here is how to do it:

1. as cricket, untar the package into /tmp:

$ cd /tmp
$ tar xvzf /home/nfs/snmp/genRtrConfig-1.4.tar.gz

2. Copy the directory /tmp/sample-config/cisco-routers and its contents to ∼/cricket-config:
$ cp -a /tmp/sample-config/cisco-routers ∼/cricket-config

3. Copy the contents of /tmp/util to ∼/cricket/util:

$ cp -a /tmp/util/* ∼/cricket/util
4. Edit ∼/cricket/util/genRtrConfig and change the first line from #!/usr/local/bin/perl
to #!/usr/bin/perl

5. Make a host table entry for the switch in /etc/hosts.

6. Change directory to ∼/cricket-config/cisco-routers and run the program genRtrConfig

like this:

$ ∼/cricket/util/genRtrConfig -C ⟨community-string⟩ --chassis \

--ciscoint -v ⟨hostname-of-switch⟩

7. Add /cisco-routers/⟨switch-hostname⟩ to Cricket’s subtree-sets file.

8. Run ∼/cricket/compile. It should now collect data. Test it using the collector.

2.6 Running the Name Service Caching Daemon

In the past, some people experienced problems unless the name service caching daemon (nscd) is running on their machines. This service is very important to reduce the load on our ldap server, and will improve the performance of your machine also, and should always be turned on. To turn on nscd:

1. First, start it now:

$ sudo service nscd start

Configuring Cricket to Monitor SNMP Objects Systems and Network Management

15

2. Now make sure that it always starts as a service when the computer boots:

     $ sudo chkconfig --level 345 nscd on

This turns the nscd service on in runlevels 3, 4 and 5.

3. Finally, verify that the service is configured correctly:

     $ chkconfig nscd --list
     nscd            0:off   1:off   2:off   3:on    4:on    5:on    6:off

If you do not see “on” for runlevels 3, 4 and 5, something is wrong (perhaps a typing mistake?)

2.7 Installing the Red Hat Updates

You will also benefit by installing the Red Hat Linux updates. These are in the di-

rectory /home/nfs/redhat-8.0/updates on ictlab, assuming ictlab:/var/ftp/pub is

mounted on /home/nfs. You can install them like this:

$ cd /home/nfs/redhat-8.0/updates/i686
$ sudo rpm -Fhv *.rpm
$ cd ../i386
$ sudo rpm -Fhv *.rpm ../noarch/*.rpm